A centrifuge rotor has a moment of inertia of 4.28×10−2 kg⋅m2 .How much energy is required to bring it from rest to 5520 rpm ?
Answer #1

We know the kinetic energy of a rotating body is, E = 1/2 I w^2

Where I is moment of inertia of the rotating body = 4.28*10^-2 kg.m^2

and w is angular velocity of the body = 5520 rpm = (5520/60) rps = 2 * 3.14* (5520/60) rad/s = 578.053 rad/s

there the energy required = 1/2 * 4.28*10^-2 * 578.053 = 12.37 joule.